OpenStudy (anonymous):

If an object moves along the y axis(marked in feet) so that its position is at time x(in seconds) is given by f(x)=88x-11x^2 A: The instantaneous velocity function v= f ' (x) B: The velocity when x=0 and x=1 seconds C: The time(s) when v=0

OpenStudy (anonymous):

A: \[v(x)=f'(x)=88-22x\]B: \[v(0)=88-22(0)=88 \]\[v(1)=88-22(1)=88-22=66\]C: \[88-22x=0\rightarrow x=4\]

OpenStudy (anonymous):

Note that velocity is in ft/sec.

OpenStudy (anonymous):

Well it must be nice to be so intelligent haha...takes you a minute to answer and I've been staring at this question for about 20 minutes with ZERO clue

OpenStudy (anonymous):

nice @genius12 :)